# CryptoMamba: Leveraging State Space Models for Accurate Bitcoin Price Prediction

### v1

@@ Abstract
Index Terms:
Bitcoin, Machine Learning, Cryptocurrency, State Space Models
I Introduction
Predicting Bitcoin [1] prices is a critical problem due to the high volatility and unpredictability of the cryptocurrency market [2]. With the growing influence of cryptocurrencies, the demand for reliable prediction models continues to rise, especially for those seeking to capitalize on market opportunities or safeguard against losses. A successful solution to this problem would benefit traders, institutions, and regulators by offering deeper insights into market behavior, enhancing decision-making, and improving the overall stability of cryptocurrency markets.
@@ I Introduction
Addressing these challenges requires models capable of effectively modeling temporal dependencies and adapting to dynamic market conditions. However, traditional statistical methods, such as ARIMA [6] and GARCH models [7], often fall short in handling complex non-linearities and sudden regime changes. In contrast, deep learning models like LSTMs (Long Short-Term Memory) [8] and Transformers [9] have shown promise in learning complex nonlinear patterns but remain limited in scalability and generalizability.
State Space Models (SSMs) [10] offer a promising alternative by modeling time series data as a combination of latent states and observed variables. These models are particularly well-suited for handling the temporal and stochastic characteristics of financial data. Recently, advances in SSMs, such as [11], have demonstrated their capability to capture long-range dependencies in sequences more effectively than traditional recurrent models. Despite their success in natural language processing [12, 13] and computer vision [14], the application of SSMs to financial time series, particularly in cryptocurrency markets, remains unexplored.
@@
  • Define and evaluate two trading algorithms, Vanilla and Smart, to assess the real-world application of predictions.
@@
Our work bridges the gap between advancements in SSMs and their practical applications in financial prediction, paving the way for future research in adaptive and robust market forecasting techniques.
II Related Work
Early attempts at Bitcoin price forecasting predominantly employed classical time series models such as ARIMA (AutoRegressive Integrated Moving Average) and GARCH (Generalized Autoregressive Conditional Heteroskedasticity), which have been widely used for financial time series prediction due to their ability to model linear relationships and volatility clustering, respectively. While ARIMA models, introduced by [6], are commonly used in financial forecasting [15, 16, 17], they assume constant variance for the data, limiting their ability to model time-varying volatility. To address this limitation, GARCH models [7] have been employed, as they excel at capturing volatility clustering, a key feature in financial markets. For instance, [18, 19] applied GARCH to model Bitcoin’s volatility, demonstrating its effectiveness for short-term forecasting. However, both ARIMA and GARCH struggle with the non-linearities and sudden regime changes that are typical of cryptocurrency markets.
@@ II Related Work
Despite the advancements in machine learning and time series models, existing approaches often struggle to effectively capture the long-range dependencies and regime shifts that characterize cryptocurrency markets. One promising approach to address the stochastic nature of Bitcoin prices is the use of State Space Models (SSMs) [11, 10]. SSMs offer a robust framework for time series analysis, particularly well-suited for capturing long-range dependencies [11], making them strong candidates for Bitcoin price prediction. Unlike traditional methods such as ARIMA and GARCH, which focus on linear relationships and short-term dependencies [19, 15], SSMs provide greater flexibility in modeling non-linearities and complex temporal interactions. Furthermore, compared to recurrent neural networks (e.g., LSTMs and GRUs), SSMs are computationally efficient, making them better suited for learning dependencies across longer time horizons. These advantages motivate the investigation of SSMs in the context of Bitcoin price forecasting.
Recently, Mamba [12], a novel variant of SSM, has gained attention for its ability to address the challenges of sequence modeling by introducing a selectivity mechanism that adapts to input data. Building on this foundation, [26] proposed S-Mamba, a model specifically designed for time series forecasting. S-Mamba employs a bidirectional Mamba layer to encode inter-variate correlations and a feed-forward network to extract temporal dependencies. S-Mamba not only outperforms state-of-the-art models in accuracy but also significantly reduces computational overhead, making it particularly effective for high-dimensional time series forecasting tasks. These advancements suggest Mamba-based models as promising candidates for financial applications like Bitcoin price forecasting.
III Background
@@
IV Methodology
CryptoMamba leverages Mamba-based SSMs to tackle the challenges of Bitcoin price prediction, offering a robust approach to capturing long-range dependencies in highly volatile financial data. This section provides an overview of the dataset utilized, the components of CryptoMamba, and the metrics employed to evaluate its performance against baseline models.
IV-A Dataset
A critical gap in Bitcoin price prediction literature is the inconsistent datasets, with studies often focusing on either early adoption periods [18] or more recent years with higher trading volumes [23]. These inconsistencies hinder the accurate assessment of the models’ generalization capabilities.To address this, we use the most recent publicly available dataset, reflecting current market trends, and evaluate CryptoMamba against baseline models to assess its effectiveness under real-world conditions.
We use historical daily Bitcoin price data from Yahoo Finance covering the period from September 17, 2018, to September 17, 2024, for our experiments. The data was partitioned into distinct intervals as shown in Table I. For our experiments, we isolate the test and the validation sets to evaluate model performance on genuinely unseen data, enabling a more robust assessment of the model’s generalization capacity.
TABLE I: Dataset splits
@@ IV-A Dataset
The data includes five main features: open, close, high, low, and volume. Since the intervals in our dataset are daily, the open and close prices represent the value of Bitcoin at the start and end of each day, which in Yahoo Finance data corresponds to the UTC time zone. The high and low prices are the maximum and minimum values recorded during each day, highlighting price volatility. Volume represents the number of Bitcoin units traded over the day, providing potential insights into market activity and sentiment that may affect price trends. Table II summarizes different features and their descriptions.
TABLE II: Dataset specifications
@@ IV-B CryptoMamba Architecture
IV-C Metric
For evaluating the forecasting accuracy of our models, we use three standard metrics: Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), and Mean Absolute Error (MAE). The formulas for these metrics are presented in Table III, where $y_{i}$ represents the actual value, $\hat{y}_{i}$ is the predicted value, and $n$ is the total number of predictions.
TABLE III: Definitions of the evaluation metrics
Metric Formula
RMSE $\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$
MAPE $\frac{100}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|$
MAE $\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
V Experiments
To evaluate the effectiveness of CryptoMamba in Bitcoin price prediction, we conduct comprehensive experiments comparing its performance with several baseline models: LSTM, Bi-LSTM, GRU, and S-Mamba. The evaluation focuses on two aspects: prediction accuracy, measured using RMSE, MAPE, and MAE, and model efficiency, quantified by the number of parameters. Additionally, we assess the impact of incorporating trading volume as a feature by conducting experiments under two setups: with and without volume. This section presents the experimental setup, results, and analysis, providing insights into the advantages of CryptoMamba over traditional and state-of-the-art baselines.
V-A Setup
To evaluate the performance of CryptoMamba, we compare it with four widely used baseline models: LSTM, Bi-LSTM, GRU, and S-Mamba. These models represent a diverse set of approaches for time-series forecasting, ranging from traditional recurrent architectures to advanced state space models.
  • LSTM (Long Short-Term Memory) [23], [22]: Configured with 3 layers and a hidden size of 100. LSTM is a popular recurrent neural network architecture and is chosen for its ability at capturing long-term dependencies in sequential data. It is widely used in financial forecasting tasks.
@@ V-A Setup
(e) S-Mamba
Figure 2: Forecasting results for all models (a) CryptoMamba, (b) LSTM, (c) Bi-LSTM, (d) GRU, and (e) S-Mamba on the training, validation, and test sets without using volume data. Non-Mamba models struggle to capture large price fluctuations, often underperforming during periods of high volatility.
V-B Evaluation
The results of our experiments, conducted with and without trading volume on the test data, are summarized in Table IV. These results show that CryptoMamba consistently outperforms all baseline models, including LSTM, Bi-LSTM, GRU, and S-Mamba, across all evaluation metrics. This demonstrates CryptoMamba’s superior ability to capture the complex dynamics of Bitcoin price movements with high generalization. Notably, CryptoMamba-v, the volume-inclusive variant, achieves the best performance with an RMSE of 1598.1, a MAPE of 0.02034, and an MAE of 1120.7. Even without volume, CryptoMamba surpasses the volume-inclusive versions of all baseline models, highlighting its robustness.
@@ V-B Evaluation
Figure 3: Forecasting results for all models (a) CryptoMamba, (b) LSTM, (c) Bi-LSTM, (d) GRU, and (e) S-Mamba on the training, validation, and test sets with volume data included as an additional feature. While volume data helps non-Mamba models slightly, they still face challenges in accurately predicting large price changes, especially during volatile periods.
Among the baselines, S-Mamba performs competitively, particularly in the volume-inclusive setup, with an RMSE of 1651.6. This highlights the strength of advanced state space models in capturing long-range dependencies. However, CryptoMamba’s tailored architecture further improves upon these results, demonstrating the benefits of its design for financial time-series forecasting. Similarly, Bi-LSTM and LSTM show significant improvements when volume data is included, whereas GRU sees minimal gains, indicating that the impact of volume depends on the model architecture.
Overall, the results confirm CryptoMamba’s effectiveness in both setups, with the inclusion of volume data consistently enhancing prediction accuracy across most models. This highlights the importance of trading volume as a feature in capturing market dynamics, particularly in highly volatile settings like Bitcoin price forecasting.
TABLE IV: Results on test data, -v shows including volume in the data
Method RMSE MAPE MAE
LSTM [23], [22] 2672.7 3.609 2094.3
Bi-LSTM [23] 2325.6 3.072 1778.8
GRU [23] 1892.4 2.385 1371.2
S-Mamba [26] 1717.4 2.248 1239.9
CryptoMamba 1713.0 2.171 1200.9
LSTM-v 2202.1 2.896 1668.9
Bi-LSTM-v 2080.2 2.738 1562.5
GRU-v 1978.0 2.526 1454.3
S-Mamba-v 1651.6 2.215 1209.7
CryptoMamba-v 1598.1 2.034 1120.7
Figures 2 and 3 illustrate the forecasting results of all models on the training, validation, and test sets, without and with volume data, respectively. While all models perform well on the training set, the non-Mamba baselines, struggle to maintain accuracy on the test set during volatile periods, showing significant divergence from actual values. S-Mamba performs better than the recurrent models but still falls short in capturing large price fluctuations. In contrast, CryptoMamba consistently tracks actual price trends, demonstrating superior generalization and robustness across both setups. The inclusion of volume data further improves the performance of most models, with CryptoMamba leveraging this additional information to achieve the most accurate predictions, particularly during periods of high volatility.
V-C Efficiency
Table V compares the number of parameters across all models, highlighting the efficiency of CryptoMamba in terms of model complexity. With only 136k parameters, CryptoMamba has the lowest parameter count among all models, significantly outperforming the much larger Bi-LSTM (569k parameters) and S-Mamba (330k parameters). Notably, LSTM and GRU, with 204k and 153k parameters, respectively, also have higher complexity than CryptoMamba. This compact architecture demonstrates CryptoMamba’s ability to effectively capture the essential patterns in the data without relying on a large number of parameters.
TABLE V: Number of parameters
Method Parameters
LSTM [23], [22] 204k
Bi-LSTM [23] 569k
GRU [23] 153k
S-Mamba [26] 330k
CryptoMamba 136k
Despite its smaller parameter size, CryptoMamba achieves state-of-the-art performance as shown in Table IV, indicating an optimal balance between efficiency and accuracy. The reduced parameter count not only lowers computational requirements, making CryptoMamba more suitable for deployment in resource-constrained environments, but also minimizes the risk of overfitting, ensuring better generalization on unseen data. Compared to S-Mamba, which has a moderately high parameter count of 330k, CryptoMamba achieves superior accuracy with less than half the complexity, further validating the benefits of its architecture for financial time-series forecasting.
VI Application in Real-World Trading
@@
In this section, we evaluate the practical utility of CryptoMamba by applying its predictions in a real-world trading scenario. Using two trading algorithms, we simulate trading with an initial balance of 100 dollars at the beginning of the validation and test period. The goal is to compute the net worth at the end of each period and compare the profits of different models.
To evaluate model predictions, we use two trading algorithms: Vanilla and Smart. Algorithm 1 describes the Vanilla Trading Algorithm, which makes simple buy or sell decisions based on the predicted and actual prices. At its core, the algorithm computes the ratio of the absolute difference between the predicted price and today’s price to today’s price. This ratio, referred to as the change ratio $d$, determines whether any trading action is required. If the change ratio is below a predefined threshold (0.01 in our case), the algorithm refrains from making any trades. This safeguard is introduced to account for transaction fees, which can render small trades unprofitable, especially when the predicted price change is minor. However, if the change ratio meets or exceeds the threshold, the algorithm takes decisive action.
@@
Algorithm 2, the Smart Trading Algorithm, introduces a risk-aware approach by considering an interval around the predicted price, determined by a risk percentage (set to 2% based on the validation MAPE of the models). This interval, defined by an upper bound ($y_{\text{max}}$) and a lower bound ($y_{\text{min}}$), accounts for prediction uncertainty and provides a range within which tomorrow’s price is likely to fall:
$y_{\text{max}}=\left(1+\frac{\text{risk}}{100}\right)y,\\ \quad y_{\text{min}}=\left(1-\frac{\text{risk}}{100}\right)y,$
where $y$ is the predicted price.
The algorithm makes decisions based on the position of today’s price ($x$) relative to this interval:
  • If $x>y_{\text{max}}$: Sell all shares, as tomorrow’s price is expected to drop significantly.
  • If $y<x\leq y_{\text{max}}$: Sell a fraction of shares, proportional to the difference between $x$ and $y$.
  • If $x<y_{\text{min}}$: Buy with all available money, as tomorrow’s price is expected to rise significantly.
  • If $y_{\text{min}}\leq x<y$: Buy a fraction of available funds, proportional to the difference between $y$ and $x$.
By dynamically adjusting trading actions based on this interval, the Smart Trading Algorithm balances risk and returns, offering a more adaptive and realistic trading strategy compared to the threshold-based Vanilla algorithm.
@@
TABLE VI: The final balance after one year of trading with \$100 in the validation and test period. Final balances that are lower than the initial investment are in Italic font.
Final Balance (\$)
@@
Model Vanilla Smart Vanilla Smart
LSTM 133.10 128.93 111.23 103.66
LSTM-v 136.93 149.95 102.09 102.06
Bi-LSTM 134.78 141.36 100.00 100.00
Bi-LSTM-v 156.98 134.41 122.40 109.96
GRU 115.12 109.06 118.46 112.64
GRU-v 153.57 125.04 104.82 106.96
S-Mamba 75.94 86.77 146.08 173.00
S-Mamba-v 77.25 91.05 182.63 203.17
CryptoMamba 113.24 106.8 175.05 184.56
CryptoMamba-v 124.09 127.12 246.58 213.20
(a) Without volume
(b) With volume
Figure 4: Net worth during trading in the test period with the vanilla trading algorithm
(a) Without volume
(b) With volume
Figure 5: Net worth during trading in the test period with the smart trading algorithm
The trading results during the test period are presented in Figures 4 and 5 for the Vanilla and Smart Trading Algorithms, respectively, with the validation and test period final net worth summarized in Table VI. In the test period, CryptoMamba consistently achieves the highest returns in both Vanilla and Smart trading setups, with CryptoMamba-v ending with \$246.58 in the Vanilla setup and \$213.20 in the Smart setup, outperforming all baselines. In the validation period, CryptoMamba-v again demonstrates strong performance, achieving \$124.09 and \$127.12 in Vanilla and Smart setups, respectively. These results underscore CryptoMamba’s ability to generalize effectively across different market conditions and adapt to varying levels of volatility.
The difference between the validation and test periods lies in the price movement dynamics: the validation period features slight, steady changes, whereas the test period exhibits large price fluctuations and volatility. Traditional baselines, such as Bi-LSTM, perform well during the validation period. However, they consistently underperform, particularly in the test period with higher volatility. This can be attributed to their lack of generalization and tendency to predict lower prices in the test period (Figures 2 and 3). In contrast, S-Mamba demonstrates strong performance in the test period, particularly under the Smart setup with \$203.17, underscoring its generalization ability and strength in capturing long-range dependencies. However, it struggles in the validation period, with final balances lower than the initial investment, suggesting it is less effective in stable market conditions.
@@
CryptoMamba exhibits superior generalization across both scenarios. Unlike the baselines, CryptoMamba performs consistently well in both steady and volatile intervals, translating accurate predictions into tangible financial gains. Its robust design allows it to adapt dynamically to market trends, achieving consistent profitability in all setups and outperforming state-of-the-art baselines in both the validation and test periods.

### v2

@@ Abstract
1 Introduction
Predicting Bitcoin [1] prices is a critical problem due to the high volatility of the cryptocurrency market [2]. With the growing influence of cryptocurrencies, the demand for reliable prediction models continues to rise, especially for those seeking to capitalize on market opportunities or safeguard against losses. A successful solution to this problem would benefit traders, institutions, and regulators by offering deeper insights into market behavior, enhancing decision-making, and improving the stability of cryptocurrency markets.
@@ 1 Introduction
Addressing these challenges requires models capable of effectively modeling temporal dependencies and adapting to dynamic market conditions. However, traditional statistical methods, such as ARIMA [6] and GARCH [7], often fall short in handling complex non-linearities and sudden regime changes. In contrast, deep learning models like LSTMs (Long Short-Term Memory) [8] and Transformers [9] show promise in learning complex nonlinear patterns but remain limited in scalability and generalizability.
State Space Models (SSMs) [10] provide a promising alternative by modeling time series data as a combination of latent states and observed variables. These models are well-suited for handling the temporal and stochastic characteristics of financial data. Recent advances in SSMs, such as [11], demonstrate their capability to capture long-range dependencies in sequences more effectively than traditional recurrent models. Despite their success in natural language processing [12, 13] and computer vision [14], the application of SSMs to financial time series, particularly in cryptocurrency markets, remains unexplored.
@@
  • Define and evaluate three trading algorithms, Vanilla, Smart, and Extended Smart, to assess the real-world application of predictions.
@@
Our work bridges the gap between advancements in SSMs and their practical financial applications, paving the way for future research in adaptive and robust market forecasting techniques.
2 Related Work
Early attempts at Bitcoin price forecasting predominantly employed classical time series models such as ARIMA (AutoRegressive Integrated Moving Average) and GARCH (Generalized Autoregressive Conditional Heteroskedasticity), widely adopted for financial time series prediction due to their ability to model linear relationships and volatility clustering, respectively. While ARIMA models, introduced by [6], are commonly used in financial forecasting [15, 16, 17], they assume constant variance for the data, limiting their ability to model time-varying volatility. To address this limitation, GARCH models [7] have been employed, as they excel at capturing volatility clustering, a key feature in financial markets. For instance, [18, 19] applied GARCH to model Bitcoin’s volatility, demonstrating its effectiveness for short-term forecasting. However, both ARIMA and GARCH struggle with the non-linearities and sudden regime changes that are typical of cryptocurrency markets.
@@ 2 Related Work
In addition to recurrent architectures, Transformer models have gained traction in time-series forecasting due to their ability to model long-range dependencies via self-attention mechanisms. iTransformer [24] adapts the Transformer architecture specifically for time-series data by introducing frequency-enhanced and hierarchical attention mechanisms, achieving state-of-the-art performance across several benchmarks. However, in the context of financial forecasting, particularly for volatile and non-stationary assets like Bitcoin, Transformer-based models often suffer from overfitting and lack robustness when trained on limited data. These shortcomings make them less suitable for high-volatility markets, motivating the use of more structured and efficient alternatives.
@@
Despite advances in machine learning and time series models, existing approaches often struggle to effectively capture the long-range dependencies and regime shifts that characterize cryptocurrency markets. One promising approach to address the stochastic nature of Bitcoin price is State Space Models (SSMs) [11, 10]. SSMs offer a robust framework for time series analysis, particularly well-suited for capturing long-range dependencies [11], making them strong candidates for Bitcoin price prediction. Unlike traditional methods such as ARIMA and GARCH, which focus on linear relationships and short-term dependencies [19, 15], SSMs provide greater flexibility in modeling non-linearities and complex temporal interactions. Furthermore, compared to recurrent neural networks (e.g., LSTMs and GRUs), SSMs are computationally efficient, enabling them to better handle dependencies across longer time horizons. These advantages motivate the investigation of SSMs in the context of Bitcoin price forecasting.
Recently, Mamba [12], a novel variant of SSM, has gained attention for introducing a selectivity mechanism that adapts model parameters to the input data. Building on this foundation, [27] proposed S-Mamba, a model specifically designed for time series forecasting. S-Mamba employs a bidirectional Mamba layer to encode inter-variate correlations and a feed-forward network to extract temporal dependencies. S-Mamba not only outperforms state-of-the-art models in accuracy but also significantly reduces computational overhead, making it particularly effective for high-dimensional time series forecasting tasks. These advancements suggest Mamba-based models as promising candidates for financial applications like Bitcoin price forecasting.
3 Background
@@
4 Methodology
CryptoMamba leverages Mamba-based SSMs to tackle the challenges of Bitcoin price prediction, offering a robust approach to capturing long-range dependencies in highly volatile financial data. This section provides an overview of the dataset utilized and the components of CryptoMamba.
4.1 Dataset
A critical gap in Bitcoin price prediction literature is the lack of unified datasets, with studies often focusing on either early adoption periods [18] or more recent years with higher trading volumes [23]. These inconsistencies hinder the accurate assessment of the models’ generalization capabilities. To address this, we use the most recent publicly available dataset, reflecting current market trends, and evaluate CryptoMamba against baseline models to assess its effectiveness under real-world conditions.
We use historical daily Bitcoin price data from Yahoo Finance covering the period from September 17, 2018, to September 17, 2024, for our experiments. The data was partitioned into distinct intervals as shown in Table 2. For our experiments, we isolate the test and the validation sets to evaluate model performance on genuinely unseen data, enabling a more robust assessment of the model’s generalization capacity.
Table 1: Dataset splits
@@ 4.1 Dataset
Table 2: Dataset specifications
@@
The data includes five main features: open, close, high, low, and volume. Since the intervals in our dataset are daily, the open and close prices represent the value of Bitcoin at the start and end of each day, which in Yahoo Finance data corresponds to the UTC time zone. The high and low prices are the maximum and minimum values recorded during each day, highlighting price volatility. Volume represents the number of Bitcoin units traded over the day, providing potential insights into market activity and sentiment that may affect price trends. Table 2 summarizes different features and their descriptions.
@@ 4.2 CryptoMamba Architecture
5 Experiments
To evaluate the effectiveness of CryptoMamba in Bitcoin price prediction, we conduct comprehensive experiments comparing its performance with several baseline models: LSTM, Bi-LSTM, GRU, iTransformer, and S-Mamba. The evaluation focuses on two aspects: prediction accuracy, measured using RMSE, MAPE, and MAE (further explained in detail in Appendix A), and model efficiency, quantified by the number of parameters. Additionally, we assess the impact of incorporating trading volume as a feature by conducting experiments under two setups: with and without volume. This section presents the experimental setup, results, and analysis, providing insights into the advantages of CryptoMamba over traditional and state-of-the-art baselines.
5.1 Setup
To evaluate the performance of CryptoMamba, we compare it with five widely used baseline models: LSTM, Bi-LSTM, GRU, iTransformer, and S-Mamba. These models represent a diverse set of approaches for time-series forecasting, ranging from traditional recurrent architectures to advanced state space models.
  • LSTM (Long Short-Term Memory) [23]: Configured with 3 layers and a hidden size of 100. LSTM is a popular recurrent neural network architecture and is chosen for its ability at capturing long-term dependencies in sequential data. It is widely used in financial forecasting tasks.
@@ 5.1 Setup
  • iTransformer[24]: Configured with d_model 128, n_heads 16, d_ff 128, and e_layers 2. iTransformer is a transformer-based model specifically designed for time-series forecasting and is included to assess the effectiveness of advanced attention-based architectures for Bitcoin price prediction.
@@ 5.2 Evaluation
(e) iTransformer
(f) S-Mamba
Figure 2: Forecasting results for all models (a) CryptoMamba, (b) LSTM, (c) Bi-LSTM, (d) GRU, (e) iTransformer, and (f) S-Mamba on the training, validation, and test sets without using volume data. Non-Mamba models struggle to capture large price fluctuations, often underperforming during periods of high volatility.
The results of our experiments, conducted with and without trading volume on the test data, are summarized in Table 3. These results show that CryptoMamba consistently outperforms all baseline models, including LSTM, Bi-LSTM, GRU, iTransformer, and S-Mamba, across all evaluation metrics. This demonstrates CryptoMamba’s superior ability to capture the complex dynamics of Bitcoin price movements with high generalization. Notably, CryptoMamba-v, the volume-inclusive variant, achieves the best performance with an RMSE of 1598.1, a MAPE of 2.034, and an MAE of 1120.7. Even without volume, CryptoMamba surpasses the volume-inclusive versions of all baseline models, highlighting its robustness.
Among the baselines, S-Mamba performs competitively, particularly in the volume-inclusive setup, with an RMSE of 1651.6. This highlights the strength of advanced state space models in capturing long-range dependencies. However, CryptoMamba’s tailored architecture further improves upon these results, demonstrating the benefits of its design for financial time-series forecasting. The iTransformer baseline performs better than traditional recurrent methods in most metrics. However, it still falls short of the performance achieved by Mamba-based models, suggesting that while attention mechanisms are effective, SSMs offer a more suitable framework for handling long-range dependencies in highly volatile financial data. Furthermore, despite having relatively low accuracy in the volume-exclusive setup, Bi-LSTM and LSTM show significant improvements when volume data is included, whereas GRU sees minimal gains, indicating that the impact of volume depends on the model architecture.
Overall, the results confirm CryptoMamba’s effectiveness in both setups, with the inclusion of volume data enhancing prediction accuracy across most models. This highlights the importance of trading volume as a feature in capturing market dynamics, particularly in highly volatile settings like Bitcoin price forecasting.
Figures 2 and 3 illustrate the forecasting results of all models on the training, validation, and test sets, without and with volume data, respectively. While all models perform well on the training set, the non-Mamba baselines, struggle to maintain accuracy on the test set during volatile periods, showing significant divergence from actual values. S-Mamba performs better than other baselines but still falls short in capturing large price fluctuations. In contrast, CryptoMamba consistently tracks actual price trends, demonstrating superior generalization and robustness across both setups. The inclusion of volume data further improves the performance of most models, with CryptoMamba leveraging this additional information to achieve the most accurate predictions, particularly during periods of high volatility.
5.3 Efficiency
Table 3 reports the number of parameters for each model alongside their regression metrics, allowing a direct comparison of predictive performance and model complexity. With only 136k parameters, CryptoMamba has the smallest footprint among all baselines, significantly outperforming much larger models like Bi-LSTM (569k) and S-Mamba (330k). Notably, even LSTM and GRU, with 204k and 153k parameters, respectively, exhibit higher complexity than CryptoMamba. This compact design enables CryptoMamba to capture essential temporal patterns without relying on a large parameter count.
@@ 5.3 Efficiency
Figure 3: Forecasting results for all models (a) CryptoMamba, (b) LSTM, (c) Bi-LSTM, (d) GRU, (e) iTransformer, and (f) S-Mamba on the training, validation, and test sets with volume data included as an additional feature. While volume data helps non-Mamba models slightly, they still face challenges in accurately predicting large price changes, especially during volatile periods.
Table 3: Results on test data, -v shows including volume in the data
Method RMSE MAPE MAE Parameters
LSTM [23] 2672.7 3.609 2094.3 204k
Bi-LSTM [23] 2325.6 3.072 1778.8 569k
GRU [23] 1892.4 2.385 1371.2 153k
iTransformer [24] 1826.9 2.460 1334.3 201k
S-Mamba [27] 1717.4 2.248 1239.9 330k
CryptoMamba 1713.0 2.171 1200.9 136k
LSTM-v 2202.1 2.896 1668.9 204k
Bi-LSTM-v 2080.2 2.738 1562.5 569k
GRU-v 1978.0 2.526 1454.3 153k
iTransformer-v 1905.9 2.540 1387.47 201k
S-Mamba-v 1651.6 2.215 1209.7 330k
CryptoMamba-v 1598.1 2.034 1120.7 136k
Despite its smaller size, CryptoMamba achieves state-of-the-art results, demonstrating an effective trade-off between efficiency and accuracy. The reduced model complexity lowers memory and computational demands, minimizing overfitting and making CryptoMamba more suitable for deployment in resource-constrained environments. For instance, compared to S-Mamba, which requires over twice the number of parameters, CryptoMamba delivers superior predictive accuracy with substantially less complexity.
In addition to its compact architecture and strong performance, CryptoMamba is highly efficient in both training and inference, making it well-suited for real-time applications. On a single NVIDIA Quadro RTX 5000 GPU, training CryptoMamba takes only 29 minutes and consumes 516 MB of GPU memory. For inference, CryptoMamba achieves a speed of 1.17 ms per sample with a batch size of 128 on the same GPU. These results further highlight CryptoMamba’s suitability for deployment in latency-sensitive and resource-constrained environments.
6 Application in Real-World Trading
@@
In this section, we evaluate the practical utility of CryptoMamba by applying its predictions in a real-world trading scenario. Using three trading algorithms, we simulate trading with an initial balance of 100 dollars at the beginning of the validation and test period. The goal is to compute the final net worth and evaluate the risk-adjusted performance of different models. To assess risk, we use the Maximum Drawdown (MDD), a widely-used metric that measures the maximum observed loss from a peak to a trough over the trading horizon (see Appendix A for details).
@@
To evaluate model predictions, we use three trading algorithms: Vanilla, Smart, and Extended Smart. Algorithm 1 describes the Vanilla Trading Algorithm, which makes simple buy or sell decisions based on the predicted and actual prices. At its core, the algorithm computes the ratio of the absolute difference between the predicted price and today’s price to today’s price. This ratio, referred to as the change ratio $d$, determines whether any trading action is required. If the change ratio is below a predefined threshold (0.01 in our case), the algorithm refrains from making any trades. This safeguard is introduced to account for transaction fees, which can render small trades unprofitable, especially when the predicted price change is minor. However, if the change ratio meets or exceeds the threshold, the algorithm takes decisive action.
Algorithm 2, the Smart Trading Algorithm, introduces a risk-aware approach by considering an interval around the predicted price, determined by a risk percentage (set to 2% based on the validation MAPE of the models). This interval, defined by an upper bound ($y_{\text{max}}$) and a lower bound ($y_{\text{min}}$), accounts for prediction uncertainty and provides a range within which tomorrow’s price is likely to fall. The algorithm makes decisions based on the position of today’s price ($x$) relative to this interval. By dynamically adjusting trading actions based on this interval, the Smart Trading Algorithm balances risk and returns, offering a more adaptive and realistic trading strategy compared to the threshold-based Vanilla algorithm.
The Extended Smart Trading Algorithm builds upon the Smart strategy by introducing support for short positions, allowing the agent to sell Bitcoin it does not own when the predicted price is expected to fall significantly. The strategy still operates based on the predicted price interval derived from the model’s risk percentage but includes a tunable cap on the maximum allowed short position (see Appendix B for algorithm details). This addition enables more flexible trading decisions, particularly under highly volatile conditions, and offers a more comprehensive evaluation of model performance.
The trading results during the test period are presented in Figure 4, with the validation and test period final net worth and MDD summarized in Table 4. In the test period, CryptoMamba consistently achieves the highest returns in all three trading strategies, with CryptoMamba-v ending with \$246.58 in the Vanilla, \$213.20 in the Smart, and \$262.78 in the Extended Smart setup, outperforming all baselines. It also consistently maintains a low MDD, especially in Extended Smart (11.09%), indicating robust risk management. This highlights CryptoMamba’s ability to translate accurate predictions into tangible financial gains while adapting well to both market rallies and downturns.
Traditional baselines like LSTM and Bi-LSTM perform relatively well in stable conditions such as the validation period but fail to generalize in the test period, where high volatility and sharp reversals occur. This can be attributed to their lack of generalization and tendency to predict lower prices in the test period (Figures 2 and 3). S-Mamba performs better in volatile settings, benefiting from its long-range modeling capacity, yet underperforms in stable intervals. The Extended Smart strategy further emphasizes the weaknesses of simpler baselines, with many models exhibiting high drawdown and final balances below the initial investment (e.g., LSTM: \$62.04, MDD: 71.46%). In contrast, CryptoMamba maintains strong performance across all setups, balancing returns and drawdowns effectively.
(a) Vanilla - no volume
(b) Smart - no volume
(c) Extended Smart - no volume
(d) Vanilla - with volume
(e) Smart - with volume
(f) Extended Smart - with volume
Figure 4: Net worth over the test period using three trading strategies, Vanilla, Smart, and Extended Smart, under both volume-exclusive (top row) and volume-inclusive (bottom row) setups. CryptoMamba consistently achieves the highest final balance across all configurations.
Table 4: The final balance and maximum drawdown (MDD) after one year of trading with \$100 in the validation and test period. Final balances that are lower than the initial investment are in Italic.
@@
Vanilla Smart Extended Smart Vanilla Smart Extended Smart
Model FB(\$) MDD(%) FB(\$) MDD(%) FB(\$) MDD(%) FB(\$) MDD(%) FB(\$) MDD(%) FB(\$) MDD(%)
LSTM 133.10 25.38 128.93 24.09 128.93 24.09 100.00 0.00 103.66 0.25 62.04 71.46
LSTM-v 136.93 25.82 149.95 25.82 149.95 25.82 100.00 0.00 102.06 0.20 70.37 60.3
Bi-LSTM 134.78 25.82 141.36 25.82 141.36 25.82 100.00 0.00 100.00 0.00 67.97 61.63
Bi-LSTM-v 156.98 25.38 134.41 25.01 136.37 25.01 98.87 12.59 109.96 9.39 91.21 71.46
GRU 115.12 19.80 109.06 19.32 111.59 19.32 100.00 0.00 112.64 3.79 117.64 21.78
GRU-v 153.57 18.70 125.04 22.31 128.07 22.31 100.00 0.00 106.96 0.88 111.96 22.85
iTransformer 100.80 28.21 114.30 20.25 103.50 24.06 151.91 14.01 151.58 13.00 128.53 28.27
iTransformer-v 100.19 24.70 116.95 20.30 115.50 21.75 241.50 17.94 206.38 16.57 207.67 37.86
S-Mamba 75.94 31.32 86.77 21.27 82.95 26.59 146.08 33.01 173.00 24.24 155.66 28.27
S-Mamba-v 77.25 35.63 91.05 21.42 78.03 34.29 182.63 27.75 203.17 16.45 222.50 18.61
CryptoMamba 113.24 24.11 106.80 22.82 102.99 22.86 175.05 23.33 184.56 16.01 204.04 17.95
CryptoMamba-v 124.09 25.38 127.12 23.40 127.47 23.40 246.58 23.48 213.20 13.37 262.78 11.09
@@
CryptoMamba exhibits superior generalization across all scenarios. Unlike the baselines, CryptoMamba performs consistently well in both steady and volatile intervals. Its robust design allows it to adapt dynamically to market trends, achieving consistent profitability in all setups and outperforming state-of-the-art baselines in both the validation and test periods.